# EUROPEAN ORGANIZATION FOR NUCLEAR RESEARCH (CERN) CERN-EP-2024-125 LHCb-PAPER-2024-007

### v1

@@ I. Kostiuk36
S. Kotriakhova
A. Kozachuk42
P. Kravchenko42
L. Kravchuk42
M. Kreps55
P. Krokovny42
W. Krupa67
W. Krzemien40
@@ A. Kozachuk42
S. Kubis77
M. Kucharczyk39
V. Kudryavtsev42
E. Kulikova42
A. Kupsc79
B. K. Kutsenko12
D. Lacarrere47
A. Lai30
A. Lampis30
D. Lancierini54
C. Landesa Gomez45
J.J. Lane1
R. Lane53
C. Langenbruch20
J. Langer18
O. Lantwin42
T. Latham55
F. Lazzari33,t
C. Lazzeroni52
R. Le Gac12
R. Lefèvre11
A. Leflat42
S. Legotin42
M. Lehuraux55
E. Lemos Cid47
O. Leroy12
T. Lesiak39
B. Leverington20
A. Li4
H. Li70
K. Li8
L. Li61
P. Li47
P.-R. Li71
Q. Li5,7
S. Li8
T. Li5,d
T. Li70
@@ H. Li70
Y. Li5
Z. Lian4
X. Liang67
S. Libralon46
C. Lin7
T. Lin56
R. Lindner47
V. Lisovskyi48
R. Litvinov30
F. L. Liu1
G. Liu70
K. Liu71
S. Liu5,7
Y. Liu57
@@ G. Liu70
Y. L. Liu60
A. Lobo Salvia44
A. Loi30
J. Lomba Castro45
T. Long54
J.H. Lopes3
A. Lopez Huertas44
S. López Soliño45
C. Lucarelli25,m
D. Lucchesi31,q
M. Lucio Martinez76
V. Lukashenko36,51
Y. Luo6
A. Lupato31
E. Luppi24,l
K. Lynch21
X.-R. Lyu7
G. M. Ma4
R. Ma7
S. Maccolini18
F. Machefert13
F. Maciuc41
B. Mack67
I. Mackay62
L. M. Mackey67
L.R. Madhan Mohan54
M. M. Madurai52
A. Maevskiy42
D. Magdalinski36
D. Maisuzenko42
@@
A. Malinin42
T. Maltsev42
G. Manca30,k
G. Mancinelli12
C. Mancuso28,13,o
@@
D. Manuzzi23
D. Marangotto28,o
J.F. Marchand10
R. Marchevski48
U. Marconi23
S. Mariani47
C. Marin Benito44
J. Marks20
A.M. Marshall53
G. Martelli32,r
G. Martellotti34
L. Martinazzoli47
M. Martinelli29,p
D. Martinez Santos45
F. Martinez Vidal46
A. Massafferri2
R. Matev47
A. Mathad47
V. Matiunin42
C. Matteuzzi67
K.R. Mattioli14
A. Mauri60
E. Maurice14
J. Mauricio44
P. Mayencourt48
M. Mazurek40
M. McCann60
L. Mcconnell21
T.H. McGrath61
N.T. McHugh58
A. McNab61
R. McNulty21
B. Meadows64
G. Meier18
D. Melnychuk40
S Meloni29,p
F. M. Meng4
M. Merk36,76
A. Merli48
L. Meyer Garcia65
D. Miao5,7
H. Miao7
M. Mikhasenko17,f
D.A. Milanes73
A. Minotti29,p
E. Minucci67
T. Miralles11
B. Mitreska18
D.S. Mitzel18
A. Modak56
A. Mödden 18
R.A. Mohammed62
R.D. Moise16
S. Mokhnenko42
T. Mombächer47
M. Monk55,1
S. Monteil11
A. Morcillo Gomez45
G. Morello26
M.J. Morello33,s
M.P. Morgenthaler20
A.B. Morris47
A.G. Morris12
R. Mountain67
H. Mu4
Z. M. Mu6
E. Muhammad55
F. Muheim57
M. Mulder75
K. Müller49
F. Muñoz-Rojas9
R. Murta60
P. Naik59
T. Nakada48
R. Nandakumar56
T. Nanut47
I. Nasteva3
M. Needham57
N. Neri28,o
S. Neubert17
N. Neufeld47
@@ I. Nasteva3
J. Nicolini18,13
D. Nicotra76
E.M. Niel48
N. Nikitin42
P. Nogarolli3
@@ E.M. Niel48
N.S. Nolte63
C. Normand53
J. Novoa Fernandez45
G. Nowak64
C. Nunez80
H. N. Nur58
A. Oblakowska-Mucha38
V. Obraztsov42
T. Oeser16
S. Okamura24,l,47
@@ A. Okhotnikov42,
O. Okhrimenko51
R. Oldeman30,k
F. Oliva57
M. Olocco18
C.J.G. Onderwater76
R.H. O’Neil57
J.M. Otalora Goicochea3
P. Owen49
A. Oyanguren46
O. Ozcelik57
K.O. Padeken17
B. Pagare55
P.R. Pais20
T. Pajero47
A. Palano22
M. Palutan26
G. Panshin42
L. Paolucci55
A. Papanestis56
M. Pappagallo22,h
L.L. Pappalardo24,l
C. Pappenheimer64
C. Parkes61
B. Passalacqua24
G. Passaleva25
D. Passaro33,s
A. Pastore22
M. Patel60
J. Patoc62
C. Patrignani23,j
A. Paul67
C.J. Pawley76
A. Pellegrino36
J. Peng5,7
M. Pepe Altarelli26
S. Perazzini23
D. Pereima42
H. Pereira Da Costa66
A. Pereiro Castro45
P. Perret11
A. Perro47
K. Petridis53
A. Petrolini27,n
J. P. Pfaller64
H. Pham67
L. Pica33
M. Piccini32
B. Pietrzyk10
G. Pietrzyk13
D. Pinci34
F. Pisani47
M. Pizzichemi29,p
V. Placinta41
M. Plo Casasus45
F. Polci15,47
M. Poli Lener26
A. Poluektov12
N. Polukhina42
I. Polyakov47
E. Polycarpo3
S. Ponce47
D. Popov7
S. Poslavskii42
K. Prasanth57
C. Prouve45
V. Pugatch51
G. Punzi33,t
S. Qasim49
W. Qian7
N. Qin4
S. Qu4
R. Quagliani47
R.I. Rabadan Trejo55
J.H. Rademacker53
M. Rama33
M. Ramírez García80
V. Ramos De Oliveira68
M. Ramos Pernas55
M.S. Rangel3
F. Ratnikov42
G. Raven37
M. Rebollo De Miguel46
F. Redi28,i
J. Reich53
F. Reiss61
Z. Ren7
P.K. Resmi62
R. Ribatti33,s
G. R. Ricart14,81
D. Riccardi33,s
S. Ricciardi56
K. Richardson63
M. Richardson-Slipper57
K. Rinnert59
P. Robbe13
G. Robertson58
E. Rodrigues59
E. Rodriguez Fernandez45
J.A. Rodriguez Lopez73
E. Rodriguez Rodriguez45
A. Rogovskiy56
D.L. Rolf47
P. Roloff47
V. Romanovskiy42
M. Romero Lamas45
A. Romero Vidal45
G. Romolini24
F. Ronchetti48
M. Rotondo26
S. R. Roy20
M.S. Rudolph67
T. Ruf47
M. Ruiz Diaz20
R.A. Ruiz Fernandez45
J. Ruiz Vidal79,aa
A. Ryzhikov42
J. Ryzka38
J. J. Saavedra-Arias9
J.J. Saborido Silva45
R. Sadek14
N. Sagidova42
D. Sahoo74
N. Sahoo52
B. Saitta30,k
M. Salomoni29,p,47
C. Sanchez Gras36
I. Sanderswood46
R. Santacesaria34
C. Santamarina Rios45
M. Santimaria26,47
L. Santoro 2
E. Santovetti35
A. Saputi24,47
D. Saranin42
G. Sarpis57
M. Sarpis61
C. Satriano34,u
A. Satta35
M. Saur6
D. Savrina42
H. Sazak16
L.G. Scantlebury Smead62
A. Scarabotto18
S. Schael16
S. Scherl59
M. Schiller58
H. Schindler47
M. Schmelling19
B. Schmidt47
S. Schmitt16
@@ H. Schmitz17,
O. Schneider48
A. Schopper47
N. Schulte18
S. Schulte48
M.H. Schune13
R. Schwemmer47
G. Schwering16
B. Sciascia26
A. Sciuccati47
S. Sellam45
A. Semennikov42
T. Senger49
M. Senghi Soares37
A. Sergi27
N. Serra49
L. Sestini31
A. Seuthe18
Y. Shang6
D.M. Shangase80
M. Shapkin42
R. S. Sharma67
I. Shchemerov42
L. Shchutska48
T. Shears59
L. Shekhtman42
Z. Shen6
S. Sheng5,7
V. Shevchenko42
B. Shi7
Q. Shi7
E.B. Shields29,p
Y. Shimizu13
E. Shmanin42
R. Shorkin42
J.D. Shupperd67
R. Silva Coutinho67
G. Simi31,q
S. Simone22,h
N. Skidmore55
T. Skwarnicki67
M.W. Slater52
J.C. Smallwood62
E. Smith63
K. Smith66
M. Smith60
A. Snoch36
L. Soares Lavra57
M.D. Sokoloff64
F.J.P. Soler58
A. Solomin42,53
A. Solovev42
I. Solovyev42
R. Song1
Y. Song48
Y. Song4
Y. S. Song6
F.L. Souza De Almeida67
B. Souza De Paula3
E. Spadaro Norella28,o
E. Spedicato23
J.G. Speer18
@@ E. Spiridenkov42,
P. Spradlin58
V. Sriskaran47
F. Stagni47
M. Stahl47
S. Stahl47
S. Stanislaus62
E.N. Stein47
O. Steinkamp49
@@ E.N. Stein47
H. Stevens18
D. Strekalina42
Y. Su7
F. Suljik62
J. Sun30
L. Sun72
Y. Sun65
@@ F. Suljik62
P.N. Swallow52
F. Swystun54
A. Szabelski40
T. Szumlak38
Y. Tan4
M.D. Tat62
A. Terentev42
F. Terzuoli33,w
F. Teubert47
E. Thomas47
D.J.D. Thompson52
H. Tilquin60
V. Tisserand11
S. T’Jampens10
M. Tobin5
L. Tomassetti24,l
G. Tonani28,o,47
X. Tong6
D. Torres Machado2
L. Toscano18
D.Y. Tou4
C. Trippl43
G. Tuci20
N. Tuning36
L.H. Uecker20
A. Ukleja38
D.J. Unverzagt20
E. Ursov42
A. Usachov37
A. Ustyuzhanin42
U. Uwer20
V. Vagnoni23
G. Valenti23
N. Valls Canudas47
H. Van Hecke66
E. van Herwijnen60
C.B. Van Hulse45,y
R. Van Laak48
M. van Veghel36
G. Vasquez49
R. Vazquez Gomez44
P. Vazquez Regueiro45
C. Vázquez Sierra45
S. Vecchi24
J.J. Velthuis53
M. Veltri25,x
A. Venkateswaran48
M. Vesterinen55
M. Vieites Diaz47
X. Vilasis-Cardona43
E. Vilella Figueras59
A. Villa23
P. Vincent15
F.C. Volle52
D. vom Bruch12
N. Voropaev42
K. Vos76
G. Vouters10,47
C. Vrahas57
J. Wagner18
J. Walsh33
E.J. Walton1,55
G. Wan6
C. Wang20
G. Wang8
J. Wang6
J. Wang5
J. Wang4
J. Wang72
M. Wang28
N. W. Wang7
R. Wang53
@@ X. Wang8,
X. Wang70
X. W. Wang60
Y. Wang6
Z. Wang13
Z. Wang4
Z. Wang28
J.A. Ward55,1
@@ X. Wang70
N.K. Watson52
D. Websdale60
Y. Wei6
J. Wendel78
B.D.C. Westhenry53
D.J. White61
M. Whitehead58
A.R. Wiederhold55
D. Wiedner18
G. Wilkinson62
M.K. Wilkinson64
M. Williams63
M.R.J. Williams57
R. Williams54
F.F. Wilson56
W. Wislicki40
M. Witek39
L. Witola20
C.P. Wong66
G. Wormser13
S.A. Wotton54
H. Wu67
J. Wu8
Y. Wu6
K. Wyllie47

### v2

@@ I. Kostiuk36â
S. Kotriakhovaâ
A. Kozachuk42â
P. Kravchenko42â
L. Kravchuk42â
M. Kreps55â
P. Krokovny42â
W. Krupa67â
W. Krzemien40â
@@ A. Kozachuk42â
S. Kubis77â
M. Kucharczyk39â
V. Kudryavtsev42â
E. Kulikova42â
A. Kupsc79â
B. K. Â Kutsenko12â
D. Lacarrere47â
A. Lai30â
A. Lampis30â
D. Lancierini54â
C. Landesa Gomez45â
J.J. Lane1â
R. Lane53â
C. Langenbruch20â
J. Langer18â
O. Lantwin42â
T. Latham55â
F. Lazzari33,tâ
C. Lazzeroni52â
R. Le Gac12â
R. Lefèvre11â
A. Leflat42â
S. Legotin42â
M. Lehuraux55â
E. Lemos Cid47â
O. Leroy12â
T. Lesiak39â
B. Leverington20â
A. Li4â
H. Li70â
K. Li8â
L. Li61â
P. Li47â
P.-R. Li71â
Q. Â Li5,7â
S. Li8â
T. Li5,dâ
T. Li70â
@@ H. Li70â
Y. Li5â
Z. Lian4â
X. Liang67â
S. Libralon46â
C. Lin7â
T. Lin56â
R. Lindner47â
V. Lisovskyi48â
R. Litvinov30â
F. L. Â Liu1â
G. Liu70â
K. Liu71â
S. Liu5,7â
Y. Liu57â
@@ G. Liu70â
Y. L. Â Liu60â
A. Lobo Salvia44â
A. Loi30â
J. Lomba Castro45â
T. Long54â
J.H. Lopes3â
A. Lopez Huertas44â
S. LĂłpez SoliĂąo45â
C. Lucarelli25,mâ
D. Lucchesi31,qâ
M. Lucio Martinez76â
V. Lukashenko36,51â
Y. Luo6â
A. Lupato31â
E. Luppi24,lâ
K. Lynch21â
X.-R. Lyu7â
G. M. Â Ma4â
R. Ma7â
S. Maccolini18â
F. Machefert13â
F. Maciuc41â
B. Â Mack67â
I. Mackay62â
L. M. Â Mackey67â
L.R. Madhan Mohan54â
M. M. Â Madurai52â
A. Maevskiy42â
D. Magdalinski36â
D. Maisuzenko42â
@@
A. Malinin42â
T. Maltsev42â
G. Manca30,kâ
G. Mancinelli12â
C. Mancuso28,13,oâ
@@
D. Manuzzi23â
D. Marangotto28,oâ
J.F. Marchand10â
R. Marchevski48â
U. Marconi23â
S. Mariani47â
C. Marin Benito44â
J. Marks20â
A.M. Marshall53â
G. Martelli32,râ
G. Martellotti34â
L. Martinazzoli47â
M. Martinelli29,pâ
D. Martinez Santos45â
F. Martinez Vidal46â
A. Massafferri2â
R. Matev47â
A. Mathad47â
V. Matiunin42â
C. Matteuzzi67â
K.R. Mattioli14â
A. Mauri60â
E. Maurice14â
J. Mauricio44â
P. Mayencourt48â
M. Mazurek40â
M. McCann60â
L. Mcconnell21â
T.H. McGrath61â
N.T. McHugh58â
A. McNab61â
R. McNulty21â
B. Meadows64â
G. Meier18â
D. Melnychuk40â
SÂ Meloni29,pâ
F. M. Â Meng4â
M. Merk36,76â
A. Merli48â
L. Meyer Garcia65â
D. Miao5,7â
H. Miao7â
M. Mikhasenko17,fâ
D.A. Milanes73â
A. Minotti29,pâ
E. Minucci67â
T. Miralles11â
B. Mitreska18â
D.S. Mitzel18â
A. Modak56â
A. MĂśdden 18â
R.A. Mohammed62â
R.D. Moise16â
S. Mokhnenko42â
T. Mombächer47â
M. Monk55,1â
S. Monteil11â
A. Morcillo Gomez45â
G. Morello26â
M.J. Morello33,sâ
M.P. Morgenthaler20â
A.B. Morris47â
A.G. Morris12â
R. Mountain67â
H. Mu4â
Z. M. Â Mu6â
E. Muhammad55â
F. Muheim57â
M. Mulder75â
K. MĂźller49â
F. MuĂąoz-Rojas9â
R. Murta60â
P. Naik59â
T. Nakada48â
R. Nandakumar56â
T. Nanut47â
I. Nasteva3â
M. Needham57â
N. Neri28,oâ
S. Neubert17â
N. Neufeld47â
@@ I. Nasteva3â
J. Nicolini18,13â
D. Nicotra76â
E.M. Niel48â
N. Nikitin42â
P. Nogarolli3â
@@ E.M. Niel48â
N.S. Nolte63â
C. Normand53â
J. Novoa Fernandez45â
G. Nowak64â
C. Nunez80â
H. N. Â Nur58â
A. Oblakowska-Mucha38â
V. Obraztsov42â
T. Oeser16â
S. Okamura24,l,47â
@@ A. Okhotnikov42,
O. Okhrimenko51â
R. Oldeman30,kâ
F. Oliva57â
M. Olocco18â
C.J.G. Onderwater76â
R.H. OâNeil57â
J.M. Otalora Goicochea3â
P. Owen49â
A. Oyanguren46â
O. Ozcelik57â
K.O. Padeken17â
B. Pagare55â
P.R. Pais20â
T. Pajero47â
A. Palano22â
M. Palutan26â
G. Panshin42â
L. Paolucci55â
A. Papanestis56â
M. Pappagallo22,hâ
L.L. Pappalardo24,lâ
C. Pappenheimer64â
C. Parkes61â
B. Passalacqua24â
G. Passaleva25â
D. Passaro33,sâ
A. Pastore22â
M. Patel60â
J. Patoc62â
C. Patrignani23,jâ
A. Â Paul67â
C.J. Pawley76â
A. Pellegrino36â
J. Â Peng5,7â
M. Pepe Altarelli26â
S. Perazzini23â
D. Pereima42â
H.  Pereira Da Costa66â
A. Pereiro Castro45â
P. Perret11â
A. Perro47â
K. Petridis53â
A. Petrolini27,nâ
J. P. Â Pfaller64â
H. Pham67â
L. Pica33â
M. Piccini32â
B. Pietrzyk10â
G. Pietrzyk13â
D. Pinci34â
F. Pisani47â
M. Pizzichemi29,pâ
V. Placinta41â
M. Plo Casasus45â
F. Polci15,47â
M. Poli Lener26â
A. Poluektov12â
N. Polukhina42â
I. Polyakov47â
E. Polycarpo3â
S. Ponce47â
D. Popov7â
S. Poslavskii42â
K. Prasanth57â
C. Prouve45â
V. Pugatch51â
G. Punzi33,tâ
S. Â Qasim49â
W. Qian7â
N. Qin4â
S. Qu4â
R. Quagliani47â
R.I. Rabadan Trejo55â
J.H. Rademacker53â
M. Rama33â
M.  RamĂrez GarcĂa80â
V. Ramos De Oliveira68â
M. Ramos Pernas55â
M.S. Rangel3â
F. Ratnikov42â
G. Raven37â
M. Rebollo De Miguel46â
F. Redi28,iâ
J. Reich53â
F. Reiss61â
Z. Ren7â
P.K. Resmi62â
R. Ribatti33,sâ
G. R. Â Ricart14,81â
D. Riccardi33,sâ
S. Ricciardi56â
K. Richardson63â
M. Richardson-Slipper57â
K. Rinnert59â
P. Robbe13â
G. Robertson58â
E. Rodrigues59â
E. Rodriguez Fernandez45â
J.A. Rodriguez Lopez73â
E. Rodriguez Rodriguez45â
A. Rogovskiy56â
D.L. Rolf47â
P. Roloff47â
V. Romanovskiy42â
M. Romero Lamas45â
A. Romero Vidal45â
G. Romolini24â
F. Ronchetti48â
M. Rotondo26â
S. R. Â Roy20â
M.S. Rudolph67â
T. Ruf47â
M. Ruiz Diaz20â
R.A. Ruiz Fernandez45â
J. Ruiz Vidal79,aaâ
A. Ryzhikov42â
J. Ryzka38â
J. J. Saavedra-Arias9â
J.J. Saborido Silva45â
R. Sadek14â
N. Sagidova42â
D. Sahoo74â
N. Sahoo52â
B. Saitta30,kâ
M. Salomoni29,p,47â
C. Sanchez Gras36â
I. Sanderswood46â
R. Santacesaria34â
C. Santamarina Rios45â
M. Santimaria26,47â
L. Santoro 2â
E. Santovetti35â
A. Saputi24,47â
D. Saranin42â
G. Sarpis57â
M. Sarpis61â
C. Satriano34,uâ
A. Satta35â
M. Saur6â
D. Savrina42â
H. Sazak16â
L.G. Scantlebury Smead62â
A. Scarabotto18â
S. Schael16â
S. Scherl59â
M. Schiller58â
H. Schindler47â
M. Schmelling19â
B. Schmidt47â
S. Schmitt16â
@@ H. Schmitz17,
O. Schneider48â
A. Schopper47â
N. Schulte18â
S. Schulte48â
M.H. Schune13â
R. Schwemmer47â
G. Schwering16â
B. Sciascia26â
A. Sciuccati47â
S. Sellam45â
A. Semennikov42â
T. Senger49â
M. Senghi Soares37â
A. Sergi27â
N. Serra49â
L. Sestini31â
A. Seuthe18â
Y. Shang6â
D.M. Shangase80â
M. Shapkin42â
R. S. Â Sharma67â
I. Shchemerov42â
L. Shchutska48â
T. Shears59â
L. Shekhtman42â
Z. Shen6â
S. Sheng5,7â
V. Shevchenko42â
B. Shi7â
Q. Shi7â
E.B. Shields29,pâ
Y. Shimizu13â
E. Shmanin42â
R. Shorkin42â
J.D. Shupperd67â
R. Silva Coutinho67â
G. Simi31,qâ
S. Simone22,hâ
N. Skidmore55â
T. Skwarnicki67â
M.W. Slater52â
J.C. Smallwood62â
E. Smith63â
K. Smith66â
M. Smith60â
A. Snoch36â
L. Soares Lavra57â
M.D. Sokoloff64â
F.J.P. Soler58â
A. Solomin42,53â
A. Solovev42â
I. Solovyev42â
R. Song1â
Y. Song48â
Y. Song4â
Y. S. Â Song6â
F.L. Souza De Almeida67â
B. Souza De Paula3â
E. Spadaro Norella28,oâ
E. Spedicato23â
J.G. Speer18â
@@ E. Spiridenkov42,
P. Spradlin58â
V. Sriskaran47â
F. Stagni47â
M. Stahl47â
S. Stahl47â
S. Stanislaus62â
E.N. Stein47â
O. Steinkamp49â
@@ E.N. Stein47â
H. Stevens18â
D. Strekalina42â
Y. Su7â
F. Suljik62â
J. Sun30â
L. Sun72â
Y. Sun65â
@@ F. Suljik62â
P.N. Swallow52â
F. Swystun54â
A. Szabelski40â
T. Szumlak38â
Y. Tan4â
M.D. Tat62â
A. Terentev42â
F. Terzuoli33,wâ
F. Teubert47â
E. Thomas47â
D.J.D. Thompson52â
H. Tilquin60â
V. Tisserand11â
S. TâJampens10â
M. Tobin5â
L. Tomassetti24,lâ
G. Tonani28,o,47â
X. Tong6â
D. Torres Machado2â
L. Toscano18â
D.Y. Tou4â
C. Trippl43â
G. Tuci20â
N. Tuning36â
L.H. Uecker20â
A. Ukleja38â
D.J. Unverzagt20â
E. Ursov42â
A. Usachov37â
A. Ustyuzhanin42â
U. Uwer20â
V. Vagnoni23â
G. Valenti23â
N. Valls Canudas47â
H. Van Hecke66â
E. van Herwijnen60â
C.B. Van Hulse45,yâ
R. Van Laak48â
M. van Veghel36â
G. Vasquez49â
R. Vazquez Gomez44â
P. Vazquez Regueiro45â
C. VĂĄzquez Sierra45â
S. Vecchi24â
J.J. Velthuis53â
M. Veltri25,xâ
A. Venkateswaran48â
M. Vesterinen55â
M. Vieites Diaz47â
X. Vilasis-Cardona43â
E. Vilella Figueras59â
A. Villa23â
P. Vincent15â
F.C. Volle52â
D. vom Bruch12â
N. Voropaev42â
K. Vos76â
G. Vouters10,47â
C. Vrahas57â
J. Wagner18â
J. Walsh33â
E.J. Walton1,55â
G. Wan6â
C. Wang20â
G. Wang8â
J. Wang6â
J. Wang5â
J. Wang4â
J. Wang72â
M. Wang28â
N. W. Â Wang7â
R. Wang53â
@@ X. Wang8,
X. Wang70â
X. W. Â Wang60â
Y. Wang6â
Z. Wang13â
Z. Wang4â
Z. Wang28â
J.A. Ward55,1â
@@ X. Wang70â
N.K. Watson52â
D. Websdale60â
Y. Wei6â
J. Wendel78â
B.D.C. Westhenry53â
D.J. White61â
M. Whitehead58â
A.R. Wiederhold55â
D. Wiedner18â
G. Wilkinson62â
M.K. Wilkinson64â
M. Williams63â
M.R.J. Williams57â
R. Williams54â
F.F. Wilson56â
W. Wislicki40â
M. Witek39â
L. Witola20â
C.P. Wong66â
G. Wormser13â
S.A. Wotton54â
H. Wu67â
J. Wu8â
Y. Wu6â
K. Wyllie47â